# Channel Capacity of Near-Field Multiuser Communications

### v1

@@ V-B Comparison with the FF Capacity
Under the FF model, the asymptotic multicast capacity for $M\rightarrow\infty$ is given in (61), shown at the bottom of this page, which yields $\mathsf{C}_{\mathsf{MC}}^{\mathsf{F}}\simeq\mathcal{O}(\log M)$.
@@
$\mathsf{F}_{\mathsf{MC}}^{\mathsf{s}}-\mathsf{F}_{\mathsf{MC}}^{\mathsf{d}}=\log_{2}\left(1+\Gamma\right)\in(0,1],$ (62)
where $\Gamma=\frac{\min_{k=1,2}r_{k}^{2}\sigma_{k}^{2}}{\max_{k=1,2}r_{k}^{2}\sigma_{k}^{2}}\in(0,1]$.
Remark 18.
The results in (62) suggest that, within the FF context, the multicast capacity is higher when the UTs are oriented in the same direction compared to when they are in different directions. This discrepancy converges to a constant not greater than one as $M$ increases.
@@ V-C Extension to Cases of $K>2$
$\mathsf{C}_{\mathsf{MC}}\leq\log_{2}\left(1+\frac{P}{K}\sum\nolimits_{k=1}^{K}{\sigma_{k}^{-2}}\left\|\mathbf{h}_{k}\right\|^{2}\right).$ (63)
@@
Please refer to Appendix -G. ∎
@@ VI Numerical Results
In this section, numerical results for the capacities of the three channels are presented. Without otherwise specification, the simulation parameter settings are defined as follows: the frequency is set as $2.4$ GHz, $d=\frac{\lambda}{2}$ m, $A=\frac{\lambda^{2}}{4\pi}$, $\frac{P_{1}}{\sigma^{2}}=\frac{P_{2}}{\sigma^{2}}=30$ dB, $\frac{P}{\sigma_{1}^{2}}=\frac{P}{\sigma_{2}^{2}}=30$ dB, $r_{1}=10$ m, and $r_{2}=5$ m. UT 1 is located in the direction $(\theta_{1},\phi_{1})=(\frac{\pi}{3},\frac{2\pi}{3})$.
VI-A CCF
Fig. 3 illustrates the values of the CCF $\rho$ when UT 2 is in various directions for both NF and FF scenarios. It can be observed that, with an array size of $65\times 65$, the CCF $\rho_{{}_{\mathsf{N}}}$ in the NF model is very close to zero regardless of UT 2’s direction, verifying that $\lim_{M_{x},{M}_{z}\rightarrow\infty}\rho_{{}_{\mathsf{N}}}\approx 0$, which demonstrates the asymptotic orthogonality of NFC for UTs in different locations. In the FF context, the value of the CCF $\rho_{{}_{\mathsf{F}}}$ is also approximately zero when UT 2 is in a distinct direction from UT 1. However, when they are located in the same direction, their channels are fully correlated, which yields $\rho_{{}_{\mathsf{F}}}=1$.
Figure 4: MAC sum-rate capacity versus $M$.
Figure 5: BC sum-rate capacity versus $M$.
@@ VI-B1 Sum-rate capacity
Fig. 5 and Fig. 5 show the sum-rate capacity of the MAC and the BC concerning the number of array elements, respectively. We can observe that under both the MAC and BC, as $M$ increases, the NF capacity and the FF capacity follow distinctly different scaling patterns. Specifically, for relatively small values of $M$, when UT 2 is in the different direction with UT 1, i.e., $(\theta_{2},\phi_{2})=(\frac{2\pi}{3},\frac{\pi}{3})$, the sum-rate capacity under the NF and FF models are similar and exhibit a linear increase with $\log M$. This similarity is attributed to the small array size, where the differences between NF and FF propagation effects are minimal and both models are accurate. However, when $M$ is sufficiently large, the disparity in channel powers and phases across the array becomes significant. In this case, the capacity under the FF model are overestimated due to the ignorance of such variations in wave propagation. As a result, the FF capacity will grow unboundedly with $M$, potentially breaking the energy-conservation laws, which is aligned with the statements in Remark 6 and 12. Conversely, as $M$ increases, the NF sum-rate capacity approach finite upper limits accurately tracked by our derived asymptotic results. This behavior is consistent with Remark 3 and 10, revealing the superior accuracy for the NF channel model.
Additionally, when UT 2 is in the same direction with UT 1, i.e., $(\theta_{2},\phi_{2})=(\frac{\pi}{3},\frac{2\pi}{3})$, for both the MAC and BC, the capacity under the FF model are lower than in the case where the UTs are in the different directions, exhibiting a smaller scaling law with $M$. These observations corroborate Remark 7 and 13. By contrast, in the NF model, the capacity for UTs in the same direction are nearly equivalent to those when they are located in different directions, surpassing the FF counterparts.
@@ VI-B2 Capacity region
In Fig. 7 and Fig. 7, the capacity regions for the MAC and the BC are presented, respectively. It is worth to nota that the capacity regions under the NF model remain almost unchanged regardless of UT 2’s direction relative to UT 1. These regions are characterized by rectangular shapes and resemble the FF capacity regions achieved when the UTs are in the different directions. On the other hand, the FF capacity regions for the UTs located in the same direction are shaped as pentagons, diverging from the rectangular regions observed under the NF model and the FF model with UTs in different directions. In particular, the rectangular capacity regions of the NF model envelop the pentagonal regions observed in the FF model for co-directional UTs. This encapsulation underscores the superior performance of NFC in managing signal correlation and interference, offering higher channel capacity for co-directional UTs.
Moreover, Fig. 9 and Fig. 9 provide insightful visualizations of how the capacity regions for the MAC and BC evolve with changes in the number of array elements, contrasting the behaviors of NF and FF models. The visualizations demonstrate that the capacity regions under the NF model stabilize to finite areas as the array size increases, verifying the correctness of Remark 5 and 11. In contrast, the capacity regions in the FF context can expand without limitation with the number of array elements, which is unreasonable under energy consideration and justifies the statements in Remark 8 and 14.
(a) $(\theta_{2},\phi_{2})=(\frac{\pi}{3},\frac{2\pi}{3})$.
(b) $(\theta_{2},\phi_{2})=(\frac{2\pi}{3},\frac{\pi}{3})$.
Figure 8: MAC capacity region versus $M$.
(a) $(\theta_{2},\phi_{2})=(\frac{\pi}{3},\frac{2\pi}{3})$.
(b) $(\theta_{2},\phi_{2})=(\frac{2\pi}{3},\frac{\pi}{3})$.
Figure 9: BC capacity region versus $M$.
@@
VI-C MC
Fig. 9 demonstrates the multicast capacity versus $M$, with different values of the AOR $\xi$. For the NF case, the musticast capacity converge to finite upper limits accurately tracked by our derived asymptotic results, which are positively correlated to the values of $\xi$. This behavior validates the conclusions in Remark 16. Additionally, contrary to the patterns observed in the MAC and BC, the FF multicast capacity for co-directional UTs exceeds that for UTs in different directions by a small gap which tend to be constant as $M$ becomes large. This distinct observation verifies Remark 18 and is attributed to the unique nature of the MC, where transmitting the same message to all UTs benefits from a high level of channel correlation.
In Fig. 10(b), we present the multicast capacity as functions of the UT 2’s distance $r_{2}$ with $r_{1}=10$ fixed at $10$ m. We can observe that, for a large number of antennas, the multicast capacity under the FF model are markedly overestimated when the UTs are located near the BS, while this overestimation diminishes as the distance from the UT to the BS increase. This is because the effect of the NF is pronounced at short distances, and will be alleviated as the UT move toward the far field. Further, the two lines under NF models are almost overlapping, with a notable exception at $r_{2}=10$ m. At this point, the scenario where UT 2 is in the same location as UT 1 exhibits a markedly higher capacity, highlighting the preference for high channel correlation in multicast settings. Regarding the FF case, the gap between the two lines initially increases with $r_{2}$, reaching a peak when UTs are equidistant from the BS ($\Gamma=1$). Beyond this point, as $r_{2}$ continues to increase, the gap decreases in accordance with $\Gamma$, as demonstrated in (62).
@@ VII Conclusion
[…]
${\mathsf{C}}_{\mathsf{MAC}}=\log_{2}\det\left({\mathbf{I}}+\sigma^{-2}{\mathbf{H}}{\mathbf{H}}^{\mathsf{H}}\right).$ (A1)
Based on the Sylvester’s theorem, (A1) can be calculated as follows:
${\mathsf{C}}_{\mathsf{MAC}}=\log_{2}\det\left({\mathbf{I}}+\sigma^{-2}{\mathbf{H}}^{\mathsf{H}}{\mathbf{H}}\right)$ (A2)
$=\log_{2}\det\!\left[\begin{matrix}1\!+\!P_{1}\sigma^{-2}\lVert{\mathbf{h}}_{1}\rVert^{2}&\!1\!+\!\sqrt{P_{1}P_{2}}\sigma^{-2}{\mathbf{h}}_{1}^{\mathsf{H}}{\mathbf{h}}_{2}\\ 1\!+\!\sqrt{P_{1}P_{2}}\sigma^{-2}{\mathbf{h}}_{2}^{\mathsf{H}}{\mathbf{h}}_{1}&\!1\!+\!P_{2}\sigma^{-2}\lVert{\mathbf{h}}_{2}\rVert^{2}\\ \end{matrix}\right].$
The results of (11) can be then obtained by calculating the above determinant of the $2\times 2$ matrix and performing simple manipulations.
@@
Based on (1) and (5), $\left\|\mathbf{h}_{k}\right\|^{2}$ under the NF model can be calculated as
$\begin{split}&G_{k}=\frac{A\Psi_{k}}{4\pi r_{k}^{2}}\sum\nolimits_{m_{x}\in\mathcal{M}_{x}}\sum\nolimits_{m_{z}\in\mathcal{M}_{z}}\\ &~{}~{}\times\big{(}(m_{x}^{2}+m_{z}^{2})\epsilon_{k}^{2}-2m_{x}\epsilon_{k}\Phi_{k}-2m_{z}\epsilon_{k}\Omega_{k}+1\big{)}^{-\frac{3}{2}}.\end{split}$ (A3)
@@
$\begin{split}G_{k}&=\frac{\xi\Psi_{k}}{4\pi}\int_{-\frac{M_{z}\epsilon_{k}}{2}}^{\frac{M_{z}\epsilon_{k}}{2}}{\int_{-\frac{M_{x}\epsilon_{k}}{2}}^{\frac{M_{x}\epsilon_{k}}{2}}}\\ &~{}~{}\times(x^{2}+z^{2}-2\Phi_{k}x-2\Omega_{k}z+1)^{-\frac{3}{2}}dxdz.\end{split}$ (A6)
We can calculate the inner integral with the aid of [25, Eq. (2.264.5)] and then the outer integral with the aid of [25, Eq. (2.284.5)], which yields the results of (14). After obtaining the channel gains, the expressions of $\rho_{{}_{\mathsf{N}}}$ in (15) can be derived by the definition of the CCF with some mathematical manipulations.
@@
$G_{k}=\frac{\xi}{4\pi}\sum\nolimits_{x\in\mathcal{X}^{\prime}_{k}}{\sum\nolimits_{z\in\mathcal{Z}^{\prime}_{k}}{\delta_{k}\left(x,z\right)}}\triangleq g(\epsilon_{k}),$ (A7)
where $\mathcal{X}^{\prime}_{k}=\left\{\frac{\epsilon_{k}}{2}\pm\Phi_{k}\right\}$, and $\mathcal{Z}^{\prime}_{k}=\left\{\frac{M\epsilon_{k}}{2}\pm\Omega_{k}\right\}$. Since $\epsilon_{k}\ll 1$, we can utilizing the first-order Taylor approximation $G_{k}\approx f\left(0\right)+f^{\prime}\left(0\right)\epsilon_{k}$, leading to the results of (19). When $M\rightarrow\infty$, we can obtain $\lim_{M\rightarrow\infty}G_{k}=\frac{\xi\epsilon_{k}\sin\phi_{k}}{2\pi\sin\theta_{k}}$ by applying L’Hôpital’s rule, which yield the results of (2).
-D Proof of Lemma 2
${\mathsf{R}}_{1}^{1\rightarrow 2}=\log_{2}\det({\mathbf{I}}+{P_{1}}{\sigma^{-2}}{\mathbf{h}}_{1}{\mathbf{h}}_{1}^{\mathsf{H}}({\mathbf{I}}+{P_{2}}{\sigma^{-2}}{\mathbf{h}}_{2}{\mathbf{h}}_{2}^{\mathsf{H}})^{-1})$
=$\overset{(a)}{=}\!\log_{2}\left(1+{P_{1}}{\sigma^{-2}}\mathbf{h}_{1}^{\mathsf{H}}\left(\mathbf{I}+{P_{2}}{\sigma^{-2}}\mathbf{h}_{2}\mathbf{h}_{2}^{\mathsf{H}}\right)^{-1}\mathbf{h}_{1}\right)$ (A8)
=$\overset{(b)}{=}\log_{2}\!\left(\!1\!+\!{P_{1}}{\sigma^{-2}}\mathbf{h}_{1}^{\mathsf{H}}\bigg{(}\mathbf{I}\!-\!\frac{{P_{2}}{\sigma^{-2}}\mathbf{h}_{2}\mathbf{h}_{2}^{\mathsf{H}}}{1\!+\!{P_{2}}{\sigma^{-2}}\left\|\mathbf{h}_{2}\right\|^{2}}\bigg{)}\mathbf{h}_{1}\right),$ (A9)
where equality $(a)$ is attained by using the Sylvester’s theorem, and $(b)$ is attained by using the Woodbury matrix identity. The results of (22a) can be obtained by performing some manipulations on (A9). After the message of UT 1 is decoded, it will be removed from the received signal, and the message of UT 2 will be decoded without interference. Consequently, the rate of UT 2 is given by
${\mathsf{R}}_{2}^{1\rightarrow 2}$ $=\log_{2}\det(\mathbf{I}+\mathbf{h}_{2}\mathbf{h}_{2}^{\mathsf{H}})\overset{\left(a\right)}{=}\log_{2}\!\left(1+P_{2}\sigma^{-2}G_{2}\right),$
where $(a)$ follows from the Sylvester’s theorem. ${\mathsf{R}}_{1}^{2\rightarrow 1}$ and ${\mathsf{R}}_{2}^{2\rightarrow 1}$ can be derived following the similar steps.
-E Proof of Proposition 1
The channel gains are easily derived from (6). On this basis, the CCF can be written as follows:
$\!\rho_{{}_{\mathsf{F}}}\!=\frac{1}{M^{2}}\left|\sum\nolimits_{m_{x}\in\mathcal{M}_{x}}{\sum\nolimits_{m_{z}\in\mathcal{M}_{z}}{\mathrm{e}^{\mathrm{j}\left(m_{x}\Delta_{\Phi}+m_{z}\Delta_{\Omega}\right)}}}\right|^{2}.$ (A10)
When $\Phi_{1}\neq\Phi_{2}$ and $\Omega_{1}\neq\Omega_{2}$, $\rho_{{}_{\mathsf{F}}}$ can be derived as follows:
$\rho_{{}_{\mathsf{F}}}\!$ =$\overset{\left(a\right)}{=}\!\frac{1}{M^{2}}\!\left|\!\frac{\big{(}\mathrm{e}^{\mathrm{j}\left(\tilde{M}_{x}\!+\!1\right)\Delta_{\Phi}}\!-\!\mathrm{e}^{-\mathrm{j}\tilde{M}_{x}\Delta_{\Phi}}\big{)}\!\big{(}\mathrm{e}^{\mathrm{j}\left(\tilde{M}_{z}\!+\!1\right)\Delta_{\Omega}}\!-\!\mathrm{e}^{-\mathrm{j}\tilde{M}_{z}\Delta_{\Omega}}\big{)}}{\left(\mathrm{e}^{\mathrm{j}\Delta_{\Phi}}-1\right)\left(\mathrm{e}^{\mathrm{j}\Delta_{\Omega}}-1\right)}\!\right|^{2}$
=$\overset{\left(b\right)}{=}\frac{16\sin^{2}\left(\frac{M_{x}\Delta_{\Phi}}{2}\right)\sin^{2}\left(\frac{M_{z}\Delta_{\Omega}}{2}\right)}{M^{2}\left(1-\cos\Delta_{\Omega}\right)^{2}\left(1-\cos\Delta_{\Phi}\right)^{2}}$
×$~{}~{}~{}\times\left|\sin\frac{\Delta_{\Phi}}{2}\sin\frac{\Delta_{\Omega}}{2}-\mathrm{j}\sin\frac{\Delta_{\Phi}}{2}\sin\frac{\Delta_{\Omega}}{2}\right|^{2},$ (A11)
where $(a)$ follows from the sum of the geometric series, and $(b)$ holds due to Euler’s formula and trigonometric identities. The final results can be obtained by calculating the magnitude. For the cases $\left(\theta_{1},\phi_{1}\right)=\left(\theta_{2},\phi_{2}\right)$, $\Phi_{1}\neq\Phi_{2},\Omega_{1}=\Omega_{2}$ and $\Phi_{1}=\Phi_{2},\Omega_{1}\neq\Omega_{2}$, (A10) can be simplified as $1$, $\frac{1}{M^{2}}\left|\sum_{m_{x}=-\tilde{M}_{x}}^{\tilde{M}_{x}}{\mathrm{e}^{\mathrm{j}m_{x}\Delta_{\Phi}}}\right|^{2}$ and $\frac{1}{M^{2}}\left|\sum_{m_{z}=-\tilde{M}_{z}}^{\tilde{M}_{z}}{\mathrm{e}^{\mathrm{j}m_{z}\Delta_{\Omega}}}\right|^{2}$, respectively, which can be calculated with similar steps to (-E).
-F Proof of Theorem 3
@@
Taking the three cases together with the substitutions of $\lVert{\mathbf{h}}_{k}\rVert^{2}=G_{k}$ and $\rho=\rho_{{}_{\mathsf{N}}}$, we can obtain the results in (56), which then lead to the multicast capacity in (57).
-G Proof of Proposition 2

### v2

@@ V-B Comparison with the FF Capacity
Under the FF model, the asymptotic multicast capacity for $M\rightarrow\infty$ is given by
$\mathsf{C}_{\mathsf{MC}}\!\simeq\!\begin{cases}\!\!\log_{2}\!\left(\frac{MPA}{4\pi}\underset{k=1,2}{\min}\,\frac{\Psi_{k}}{r_{k}^{2}\sigma_{k}^{2}}\right)\!\triangleq\!\mathsf{F}_{\mathsf{MC}}^{\mathsf{s}}&\!\left(\theta_{1},\phi_{1}\right)\!=\!\left(\theta_{2},\phi_{2}\right)\\ \!\!\log_{2}\!\left(\frac{MPA/(4\pi)}{\Psi_{1}^{-1}r_{1}^{2}\sigma_{1}^{2}+\Psi_{2}^{-1}r_{2}^{2}\sigma_{2}^{2}}\right)\!\triangleq\!\mathsf{F}_{\mathsf{MC}}^{\mathsf{d}}&\!\left(\theta_{1},\phi_{1}\right)\!\neq\!\left(\theta_{2},\phi_{2}\right)\\ \end{cases}\!,$ (85)
which yields $\mathsf{C}_{\mathsf{MC}}\simeq\mathcal{O}(\log M)$.
@@
$\mathsf{F}_{\mathsf{MC}}^{\mathsf{s}}-\mathsf{F}_{\mathsf{MC}}^{\mathsf{d}}=\log_{2}\left(1+\Upsilon\right)\in(0,1],$ (86)
where $\Upsilon=\frac{\min_{k=1,2}r_{k}^{2}\sigma_{k}^{2}}{\max_{k=1,2}r_{k}^{2}\sigma_{k}^{2}}\in(0,1]$. The above results suggest that, within the FF context, the multicast capacity is higher when the UTs are oriented in the same direction compared to when they are in different directions, while the discrepancy converges to a constant not greater than one as $M$ increases.
@@ V-C Extension to Cases of $K>2$
$\mathsf{C}_{\mathsf{MC}}\leq\log_{2}\left(1+\frac{P}{K}\sum\nolimits_{k=1}^{K}{\sigma_{k}^{-2}}\mathsf{g}_{k}^{\mathsf{n}}\right).$ (87)
@@
Please refer to Appendix -E. ∎
@@ VI Numerical Results
In this section, numerical results for the capacities of the three channels are presented. Without otherwise specification, the simulation parameter settings are defined as follows: the frequency is set as $2.4$ GHz, $d=\frac{\lambda}{2}$ m, $A=\frac{\lambda^{2}}{4\pi}$, $\gamma_{1}=\gamma_{2}=30$ dB, $\frac{P}{\sigma_{1}^{2}}=\frac{P}{\sigma_{2}^{2}}=30$ dB, $r_{1}=10$ m, and $r_{2}=5$ m. UT 1 is located in the direction $(\theta_{1},\phi_{1})=(\frac{\pi}{3},\frac{2\pi}{3})$; and UT 2 is located in the same direction or the different direction, $(\theta_{2},\phi_{2})=(\frac{2\pi}{3},\frac{\pi}{3})$, with respect to UT 1, denoted as “SD” and “DD” in the figures, respectively.
VI-A Channel Correlation Factor
Fig. 6 illustrates the values of the CCF $\rho$ when UT 2 is in various directions for both NF and FF scenarios. It can be observed that, with an array size of $65\times 65$, the CCF $\rho_{\mathsf{n}}$ in the NF model is very close to zero regardless of UT 2’s direction, verifying that $\lim_{M_{x},{M}_{z}\rightarrow\infty}\rho_{\mathsf{n}}\approx 0$, which demonstrates the asymptotic orthogonality of NFC for UTs in different locations. In the FF context, the value of the CCF $\rho_{{}_{\mathsf{F}}}$ is also approximately zero when UT 2 is in a distinct direction from UT 1. However, when they are located in the same direction, their channels are fully correlated, which yields $\rho_{{}_{\mathsf{F}}}=1$.
(a) NF MAC sum-rate capacity with respect to $(\Delta\theta,\Delta\phi)$.
(b) FF MAC sum-rate capacity with respect to $(\Delta\theta,\Delta\phi)$.
Figure 9: Illustration of impacts of small perturbations in angle difference for co-directional UTs on capacity, with $M_{x}=M_{z}=45$.
@@ VI-B1 Sum-Rate Capacity
Fig. 8 and Fig. 8 show the sum-rate capacity of the MAC and the BC concerning the number of array elements, respectively. We can observe that under both the MAC and BC, as $M$ increases, the NF capacity and the FF capacity follow distinctly different scaling patterns. Specifically, for relatively small values of $M$, when UT 2 is in the different direction with UT 1, the sum-rate capacity under the NF and FF models are similar and exhibit a linear increase with $\log M$. This similarity is attributed to the small array size, where the differences between NF and FF propagation effects are minimal and both models are accurate. However, when $M$ is sufficiently large, the disparity in channel powers and phases across the array becomes significant. In this case, the capacity under the FF model is overestimated due to the ignorance of such variations in wave propagation. As a result, the FF capacity will grow unboundedly with $M$, potentially breaking the energy-conservation laws, which is aligned with the statements in Remark 4 and 8. Conversely, as $M$ increases, the NF sum-rate capacity approach finite upper limits accurately tracked by our derived asymptotic results. This behavior is consistent with Remark 1 and 6, revealing the superior accuracy for the NF channel model.
Additionally, when UT 2 is in the same direction with UT 1, for both the MAC and BC, the capacity under the FF model are lower than in the case where the UTs are in the different directions, exhibiting a smaller scaling law with $M$. These observations corroborate our previous analysis. By contrast, in the NF model, the capacity for UTs in the same direction are nearly equivalent to those when they are located in different directions, surpassing the FF counterparts.
While we find that the capacity of the MAC and BC with co-directional UTs can be improved under NFC compared to the FF case, it is nearly impossible for UTs to be positioned in the exact same direction in real-word scenarios, typically only similarly. Therefore, it would be crucial to explore how small perturbations in the angle difference between the UTs affect capacity. By letting $(\theta_{2},\phi_{2})=(\theta_{1}+\Delta\theta,\phi_{1}+\Delta\phi)$ with $(\Delta\theta,\Delta\phi)$ denoting the small angle perturbations, Fig. 9 illustrates how MAC capacity fluctuates with changes in $(\Delta\theta,\Delta\phi)$, noting that similar outcomes are expected for the BC, which are not shown for for brevity. It is clear that the NF capacity exceeds the FF capacity when the angle perturbations $\Delta\theta$ and $\Delta\phi$ are very small. More importantly, as the perturbations increase, the NF capacity grows more rapidly than the FF capacity, meaning that the NF capacity is more sensitive to the angle perturbations. In particular, compared to cases where the UTs are in the same direction, even slight angle differences can significantly enhance capacity, underscoring NFC’s better angular resolution, which confirms the superior performance of NFC for UTs in the same direction or with small angle perturbations.
@@ VI-B2 Capacity Region
In Fig. 11 and Fig. 11, the capacity regions for the MAC and the BC are presented, respectively. It is worth to note that the capacity regions under the NF model remain almost unchanged regardless of UT 2’s direction relative to UT 1. These regions are characterized by rectangular shapes and resemble the FF capacity regions achieved when the UTs are in the different directions. On the other hand, the FF capacity regions for the UTs located in the same direction are shaped as pentagons, diverging from the rectangular regions observed under the NF model and the FF model with UTs in different directions. In particular, the rectangular capacity regions of the NF model envelop the pentagonal regions observed in the FF model for co-directional UTs. This encapsulation underscores the superior performance of NFC in managing signal correlation and interference, offering higher channel capacity for co-directional UTs.
(a) MAC capacity region (SD).
(b) MAC capacity region (DD).
(c) BC capacity region (SD).
(d) BC capacity region (DD).
Figure 12: MAC channel capacity and BC channel capacity versus the antenna number $M$.
Moreover, Fig. 12 provides insightful visualizations of how the capacity regions for the MAC and BC evolve with changes in the number of array elements, contrasting the behaviors of NF and FF models. The visualizations demonstrate that the capacity regions under the NF model stabilize to finite areas as the array size increases, verifying the correctness of Remark 3 and 7. In contrast, the capacity regions in the FF context can expand without limitation with the number of array elements, which is unreasonable under energy consideration.
@@
VI-C Multicast Channel
Fig. 13(a) demonstrates the multicast capacity versus $M$, with different values of the AOR $\xi$. For the NF case, the musticast capacity converge to finite upper limits accurately tracked by our derived asymptotic results, which are positively correlated to the values of $\xi$. This behavior validates the conclusions in Remark 10. Additionally, contrary to the patterns observed in the MAC and BC, the FF multicast capacity for co-directional UTs exceeds that for UTs in different directions by a small gap which tend to be constant as $M$ becomes large, verifying our theoretical analysis. This distinct observation is attributed to the unique nature of the MC, where transmitting the same message to all UTs benefits from a high level of channel correlation.
In Fig. 13(b), we present the multicast capacity as a function of the UT 2’s distance $r_{2}$ with $r_{1}$ fixed at $10$ m. We observe that when the UTs are located near the BS, the multicast capacity under the FF model is significantly overestimated. However, this overestimation diminishes as the distance from the UT to the BS increases. This occurs because the NF effect is more pronounced at short distances and becomes less significant as the UT moves toward the far field. Further, the two lines under NF models are almost overlapping, with a notable exception at $r_{2}=10$ m. At this point, the scenario where UT 2 is in the same location as UT 1 exhibits a markedly higher capacity, highlighting the preference for high channel correlation in multicast settings. Regarding the FF case, the gap between the two lines initially increases with $r_{2}$, reaching a peak when UTs are equidistant from the BS ($\Upsilon=1$). Beyond this point, as $r_{2}$ continues to increase, the gap decreases in accordance with $\Upsilon$, as demonstrated in (86).
@@ VII Conclusion
${\mathsf{R}}_{1}^{1\rightarrow 2}$ $=\log_{2}\det({\mathbf{I}}_{M}+{\gamma_{1}}{\mathbf{h}}_{1}{\mathbf{h}}_{1}^{\mathsf{H}}({\mathbf{I}}_{M}+{\gamma_{2}}{\mathbf{h}}_{2}{\mathbf{h}}_{2}^{\mathsf{H}})^{-1})$
=$\overset{(a)}{=}\log_{2}\left(1+{\gamma_{1}}\mathbf{h}_{1}^{\mathsf{H}}\left(\mathbf{I}_{M}+{\gamma_{2}}\mathbf{h}_{2}\mathbf{h}_{2}^{\mathsf{H}}\right)^{-1}\mathbf{h}_{1}\right)$
=$\overset{(b)}{=}\log_{2}\bigg{(}1+{\gamma_{1}}\mathbf{h}_{1}^{\mathsf{H}}\bigg{(}\mathbf{I}_{M}-\frac{{\gamma_{2}}\mathbf{h}_{2}\mathbf{h}_{2}^{\mathsf{H}}}{1+{\gamma_{2}}\left\|\mathbf{h}_{2}\right\|^{2}}\bigg{)}\mathbf{h}_{1}\bigg{)},$ (A1)
where equality $(a)$ is attained by using the Sylvester’s identity, and $(b)$ is attained by using the Woodbury matrix identity. The results of (11a) can be obtained by performing some manipulations on (A1). After the message of UT 1 is decoded, it will be removed from the received signal, and the message of UT 2 will be decoded without interference. Consequently, the rate of UT 2 is given by
${\mathsf{R}}_{2}^{1\rightarrow 2}=\log_{2}\det(\mathbf{I}_{M}+\gamma_{2}\mathbf{h}_{2}\mathbf{h}_{2}^{\mathsf{H}})\overset{\left(a\right)}{=}\log_{2}\big{(}1+\gamma_{2}\left\|\mathbf{h}_{2}\right\|^{2}\big{)},$ (A2)
where $(a)$ follows from the Sylvester’s identity. ${\mathsf{R}}_{1}^{2\rightarrow 1}$ and ${\mathsf{R}}_{2}^{2\rightarrow 1}$ can be derived following the similar steps.
@@
Based on (1) and (5), the channel gain under the NF model is calculated as
$\begin{split}&\mathsf{g}_{k}^{\mathsf{n}}=\frac{A\Psi_{k}}{4\pi r_{k}^{2}}\sum\nolimits_{m_{x}\in\mathcal{M}_{x}}\sum\nolimits_{m_{z}\in\mathcal{M}_{z}}\\ &~{}~{}\times\big{(}(m_{x}^{2}+m_{z}^{2})\epsilon_{k}^{2}-2m_{x}\epsilon_{k}\Phi_{k}-2m_{z}\epsilon_{k}\Omega_{k}+1\big{)}^{-\frac{3}{2}}.\end{split}$ (A3)
@@
[…]
We can calculate the inner integral with the aid of [34, Eq. (2.264.5)] and then the outer integral with the aid of [34, Eq. (2.284)], which yields the results of (15).
Following similar steps to obtain (A6), the CCF can be written as follows:
$\rho$ $=\prod_{k=1}^{2}{\frac{A\Psi_{k}}{4\pi r_{k}^{2}\mathsf{g}^{\mathsf{n}}_{k}}}\left|\sum_{m_{x},m_{z}}{f_{1}\left(m_{x}\epsilon_{1},m_{z}\epsilon_{1}\right)g_{2}\left(m_{x}\epsilon_{1},m_{z}\epsilon_{1}\right)}\right|^{2}$
$=\prod_{k=1}^{2}{\frac{A\Psi_{k}}{4\pi r_{k}^{2}\mathsf{g}^{\mathsf{n}}_{k}\epsilon_{1}^{2}}}\left|\int_{\frac{-M_{z}\epsilon_{1}}{2}}^{\frac{M_{z}\epsilon_{1}}{2}}{\int_{\frac{-M_{x}\epsilon_{1}}{2}}^{\frac{M_{x}\epsilon_{1}}{2}}{f_{1}\left(x,z\right)f_{2}\left(x,z\right)\mathrm{d}x}\mathrm{d}z}\right|^{2}.$ (A7)
The above integrals can be numerically evaluated by applying the Chebyshev-Gauss quadrature rule, i.e., $\int_{-1}^{1}{\frac{f\left(x\right)}{\sqrt{1-x^{2}}}\mathrm{d}}x\approx\sum_{t=1}^{T}{f\left(x_{j}\right)}$ with $x_{t}=\cos\left(\frac{2t-1}{2T}\pi\right)$, resulting in the expressions of (16).
@@
$\begin{split}\mathsf{g}_{k}^{\mathsf{n}}=&\frac{\xi}{4\pi}\sum\nolimits_{x\in\left\{\frac{\epsilon_{k}}{2}\pm\Phi_{k}\right\}}\sum\nolimits_{z\in\left\{\frac{M\epsilon_{k}}{2}\pm\Omega_{k}\right\}}\\ &\times{\arctan\bigg{(}\frac{xz}{\Psi_{k}\sqrt{\Psi_{k}^{2}+x^{2}+z^{2}}}\bigg{)}}\triangleq g(\epsilon_{k}).\end{split}$ (A8)
Since $\epsilon_{k}\ll 1$, we can utilizing the first-order Taylor approximation $\mathsf{g}_{k}^{\mathsf{n}}\approx g\left(0\right)+g^{\prime}\left(0\right)\epsilon_{k}$, leading to the results of (22). When $M\rightarrow\infty$, we can obtain $\lim_{M\rightarrow\infty}\mathsf{g}_{k}^{\mathsf{n}}=\frac{\xi\epsilon_{k}\sin\phi_{k}}{2\pi\sin\theta_{k}}$ by applying L’Hôpital’s rule, which yield the results of (23).
-D Proof of Theorem 3
@@
Taking the three cases together with the substitutions of $\lVert{\mathbf{h}}_{k}\rVert^{2}=\mathsf{g}_{k}^{\mathsf{n}}$ and $\rho=\rho_{\mathsf{n}}$, we can obtain the results in (80), which then lead to the multicast capacity in (81).
-E Proof of Lemma 3